teressa24:

Mike and his friends bought cheese wafers for $2 per packet and chocolate wafers for $1 per packet at a carnival. They spent a total of $25 to buy a total of 20 packets of wafers of the two varieties. Part A: Write a system of equations that can be solved to find the number of packets of cheese wafers and the number of packets of chocolate wafers that Mike and his friends bought at the carnival. Define the variables used in the equations. (5 points) Part B: How many packets of chocolate wafers and cheese wafers did they buy? Explain how you got the answer and why you selected a particular method to get the answer. (5 points)

snowflake0531:

Do you know how you would set up the equation?

teressa24:

no

snowflake0531:

Okay, so we first look at the first sentence cheese wafers for $2 per packet and chocolate wafers for $1 per packet which means 2x + y = something, assuming that x is cheese wafers and y is chocolate wafers They spent a total of $25 to buy a total of 20 packets of wafers of the two varieties. This finishes the above expression up 2x+y=25 But then, they bought 25 in total, so x+y=20

snowflake0531:

Do you know any of the methods to solve for x and y?

snowflake0531:

aka the substitution method, elimination method, or graphing. Graphing would be easiest?

teressa24:

no i dont

snowflake0531:

Then I recommend graphing the pair of linear equations for now

snowflake0531:

And about the reason, you can just write it's easier

snowflake0531:

https://www.desmos.com/calculator/ygbbpktsqf there's the graph

snowflake0531:

you can see that it intersects at (5,15), where x is 5, y is 15 This means that he bought 5 cheese flavored, and 15 chocolate flavored